# references.bib

### v1

@@ 5.3. Case study 2: double vortex
Fig. 4 shows the centred scheme adequately resolves turbulent flow structures over long periods of time with minimal grid scale noise. Similar to other studies that consider vortex pair interactions [Giorgetta2009, Reich2006, Mcrae2013, Staniforth2006, Eldred2019], we observe the formation of two distinct vortices (Fig. 4(a)) which collide as a mature turbulent state is reached (Fig. 4(f)). The ability of the centred scheme to resolve such complex features is further evidence that the new linearised Jacobian 31 provides robust convergence with respect to large variations from the initial mean flow.
@@
For the centred scheme, Fig. 5(a,b) shows conservation of energy and mass to machine precision. Similar to Section 5.2, Fig. 5(c) shows a drift in entropy conservation from machine precision to $\mathcal{O}(10^{-8})$ as the simulation progresses even though $\partial\mathcal{S}_{h}/\partial t$ is machine zero (Fig. 5(d)). Thus, we conclude entropy is not conserved point-wise in time.
Figure 5. Conservation errors for the double vortex case study. (a,b,c) show normalised values of energy, mass and entropy, and (d) illustrates the change in entropy due to forcing terms 21. Each figure compares the centred (blue) and constrained schemes (orange). The vertical axis is logarithmic and the horizontal axis shows unit-less time. Parameters: $n_{s}=64$ spatial elements, $p=1$ spatial finite elements, $\mathrm{CFL}=0.2$.
To achieve exact entropy conservation, we use the constrained formulation in 30, for which entropy conservation is explicitly enforced using a Lagrange multiplier. Conservation errors in energy, mass and $\partial\mathcal{S}_{h}/\partial t$ are are comparable between the centred and constrained formulations (Fig. 5(a,b,d)), and Fig. 5(c) shows entropy is conserved within the tolerance of the non-linear solver. In situations where exact entropy conservation is required, we advise use of the constrained formulation.
@@ 6. Conclusion and future work
A consequence of the chosen Poisson time integrator [Cohen2011] is that cubic Casimirs are not preserved for non-canonical Hamiltonian systems. Thus, discrete entropy is not conserved point-wise in time. To address this, we propose a constrained formulation using Lagrange multipliers which numerical experiments confirm corrects small temporal losses in exact entropy conservation. While Poisson time integrators that preserve higher order Casimirs for canonical Hamiltonian systems have recently been developed [Modin2020], the extension to non-canonical Hamiltonian systems is an open problem.
@@ Acknowledgements
This research is supported by the Commonwealth of Australia as represented by the Defence Science and Technology Group of the Department of Defence. This research is also funded by the Australian Government through the Australian Research Council (project numbers DP210103092 and DP220103160). This work is supported by computational resources provided by the Australian Government through NCI under the National Computational Merit Allocation Scheme (NCMAS), ANU Merit Allocation Scheme, and the Monash-NCI partnership program.

### v2

@@ 5.3. Case study 2: double vortex
Fig. 4 shows the centred scheme adequately resolves turbulent flow structures over long periods of time with minimal grid scale noise. Similar to other studies that consider vortex pair interactions [Giorgetta2009, Reich2006, Mcrae2013, Staniforth2006, Eldred2019], we observe the formation of two distinct vortices (Fig. 4(a)) which collide as a mature turbulent state is reached (Fig. 4(f)). The ability of the centred scheme to resolve such complex features is further evidence that the new linearised Jacobian 33 provides robust convergence with respect to large variations from the initial mean flow.
@@
For the centred scheme, Fig. 5(a,b) shows conservation of energy and mass to machine precision. Similar to Section 5.2, Fig. 5(c) shows a drift in entropy conservation from machine precision to $\mathcal{O}(10^{-8})$ as the simulation progresses even though $\partial\mathcal{S}_{h}/\partial t$ is machine zero (Fig. 5(d)). Thus, we conclude entropy is not conserved pointwise in time.
Figure 5. Conservation errors for the double vortex case study. (a,b,c) show normalised values of energy, mass and entropy, and (d) illustrates the change in entropy due to forcing terms 23. Each figure compares the centred (blue) and constrained schemes (orange). The vertical axis is logarithmic and the horizontal axis shows unit-less time. Parameters: $n_{s}=64$ spatial elements, $p=1$ spatial finite elements, $\mathrm{CFL}=0.2$.
To achieve exact entropy conservation, we use the constrained formulation in 32, for which entropy conservation is explicitly enforced using a Lagrange multiplier. Conservation errors in energy, mass and $\partial\mathcal{S}_{h}/\partial t$ are comparable between the centred and constrained formulations (Fig. 5(a,b,d)), and Fig. 5(c) shows entropy is conserved within the tolerance of the non-linear solver. In our experience, solving 32 only requires a few more non-linear iterations relative to the unconstrained formulation. Thus, in situations where exact entropy conservation is required, we advise use of the constrained formulation.
@@ 6. Conclusion and future work
A consequence of the chosen Poisson time integrator [Cohen2011] is that cubic Casimirs are not preserved for non-canonical Hamiltonian systems. Thus, discrete entropy is not conserved pointwise in time. To address this, we propose a constrained formulation using Lagrange multipliers which numerical experiments confirm corrects small temporal losses in exact entropy conservation. The constrained formulation is recommended in situations where exact entropy conservation is required. However, the unconstrained upwinded scheme, for which entropy is monotonically damped, is generally preferred to smooth spurious oscillations that arise in the absence of stabilisation. From the analysis and results presented in this study, we have identified the need to develop a Poisson time integrator that preserves higher order Casimirs for Poisson systems in general. While Poisson time integrators that preserve higher order Casimirs for canonical Hamiltonian systems have recently been developed [Modin2020], the extension to non-canonical Hamiltonian systems is an open problem.
@@ Acknowledgements
This research is supported by the Commonwealth of Australia as represented by the Defence Science and Technology Group of the Department of Defence. This research is also funded by the Australian Government through the Australian Research Council (project numbers DP210103092 and DP220103160). This work is supported by computational resources provided by the Australian Government through NCI under the National Computational Merit Allocation Scheme (NCMAS), ANU Merit Allocation Scheme, and the Monash-NCI partnership program. The authors acknowledge the two anonymous reviewers, whose comments helped to improve the quality of this article.